OpenStudy (anonymous):

The graph below shows a system of equations:

OpenStudy (anonymous):

The x-coordinate of the solution to the system of equations is __________.

OpenStudy (help!!!!):

Do have any choices?

OpenStudy (help!!!!):

Well no points in this region will satisfy y=-x+5 or y=x-1

OpenStudy (anonymous):

answer choices: x = -1 x = 0 x = 1 x = 3

OpenStudy (jdoe0001):

ahemm the solution for a system of LINEar equations is at the point where both LINEs meet or intersect

OpenStudy (help!!!!):

so x=3

OpenStudy (jdoe0001):

so... where do they meet?

OpenStudy (anonymous):

@jdoe0001 is that the right answer? because i just realized that there werent any answer choices given for that question XD